# Slow is Fast! Dissecting Ethereum’s Slow Liquidity Drain Scams

### v2

@@ 2 ML_approach (SLID_pools, DEXData) :
The first function (LP_Age_Analysis) determines the longevity and activity status of each pool. It retrieves transactions associated with each pool and calculates the pool’s age by comparing the earliest and latest transaction timestamps. This data helps assess how long each pool has been operational. The function also tracks whether pools are still active by checking for recent transactions.
In the second function (LP_Profit_Analysis), the focus shifts to the financial activities of the pool owners. This part of the algorithm targets transactions where the pool owners are selling or withdrawing assets. By analyzing these transactions, the function calculates the total volume of assets moved, providing insights into the profit-taking behaviors of pool owners.
The final function (LP_Trend_Analysis) examines the engagement of other users within the pools, excluding the owners. It analyzes transactions made by users to understand the general activity level and trends within each pool. This function helps to highlight user participation and can indicate the popularity of the pool based on user activity.
We further present pseudocode for our enhance method by using machine learning (Algorithm 1).
@@ 9 LP_Age_Analysis (SLID_pools, DEXData) :
12             /* get all transaction of the pool */
[…]
14             /* pool age calculation */
[…]
@@
[…]
@@ 23 LP_Profit_Analysis (SLID_pools, DEXData) :
[…]
@@ 38 LP_Trend_Analysis (SLID_pools, DEXData) :
[…]
@@
Ethical Considerations
Privacy considerations. We focused our research on the behavior pattern of the owner that creates SLID liquidity pool and their activities. All examined data are publicly available by the Ethereum network, with all addresses anonymized by only keeping the first/last 5 characters of each, and generated mapping values for similarities determination (of determining transactions involving the same tokens/pools/users) in analysis tasks. We have no intention of deanonymizing any of the involved addresses for privacy reasons.
Respect for persons. We prioritized user privacy throughout this research. All data analyzed was sourced from publicly available blockchain records, and no personally identifiable information (PII) was collected or processed. While blockchain data is inherently pseudonymous (also claimed before), we ensured that our findings respected user anonymity. Additionally, any insights derived from community-reported rug pulls were aggregated to avoid targeting individual users.
Open Science
To promote transparency and reproducibility in research, we will share relevant details, datasets, and code used within this study available once the paper is fully accepted/finalized.
Table 4: Summerized features for datasets.
@@
LP $Address_{pool}$ Ethereum address of LPs $Address_{base}$ Ethereum address of base token in LPs
@@
$DEX$ DEX where the liquidity pool deployed $Name$ Paired token’s name
DEX activities $Block$ Block number that the activity occurred $Timestamp$ Timestamp that the activity occured
$Hash$ Transaction hash of the activity $Category$ Activities (buy/sell/deposit/withdraw)
@@
Honeypot $Tax_{buy}$ Buy tax of the token $Tax_{modify}$ Owner’s capability to change buy/sell
$Tax_{sell}$ Sell tax of the token $Buyable$ User’s capability to buy tokens
$SellAll$ User’s capability to sell all tokens $BalanceChange$ Owner’s capablity to change user’s balance
$TradingCooldown$ Time limit for tokens between activities $TradingPausable$ Owner’s capbility of pausing trading
$AntiWhale$ token activities with high amounts
@@
OAF $Owner_{dep}$ Total owner deposit activity $Owner_{buy}$ Total owner buy activity
@@
UAF $User_{dep}$ Total users deposit activity $User_{with}$ Total users withdraw activity
@@
PF $Owner_{i}$ Total owner investment (USD) $Owner_{u}$ Total owner unrealized return (USD)
@@
$ROwner_{roi}$ The ratio of ($Owner_{r}$--$Owner_{i}$) and $Owner_{i}$ $ROwner_{uonr}$ The ratio of $Owner_{u}$ and $Owner_{r}$
@@
LPF $Age$ LP age ($=d$ if the pool still active) $IsAlive$ Whether the pool is still active

### v3

@@ 2 ML_approach (SLID_pools, DEXData) :
The first function (LP_Age_Analysis) determines the longevity and activity status of each pool. It retrieves transactions associated with each LP and calculates the pool’s age by comparing the earliest and latest transaction timestamps. This data helps assess how long each LP has been operational. The function also tracks whether pools are still active by checking for recent transactions.
@@ 9 LP_Age_Analysis (SLID_pools, DEXData) :
12     /* get all transaction of the LP */
13     pool_age $\leftarrow$ $max(\texttt{pool\_txs.timestamp})-min(\texttt{pool\_txs.timestamp})$
14     /* LP age calculation */
15     ages[pool_age] $\leftarrow$ $\texttt{ages[pool\_age]}+1$
@@
17     alive[pool_age] $\leftarrow$ $\texttt{ages[pool\_age]}+1$ if max(pool_txs.timestamp) $\geq$ now
@@ 23 LP_Profit_Analysis (SLID_pools, DEXData) :
31       count[transaction.timestamp] $\leftarrow$ $\texttt{count[transaction.timestamp]}+1$
@@ 38 LP_Trend_Analysis (SLID_pools, DEXData) :
46       count[transaction.timestamp] $\leftarrow$ $\texttt{count[transaction.timestamp]}+1$
@@
In the second function (LP_Profit_Analysis), the focus shifts to the financial activities of the LP owners. This part of the algorithm targets transactions where the LP owners are selling or withdrawing assets. By analyzing these transactions, the function calculates the total volume of assets moved, providing insights into the profit-taking behaviors of LP owners.
The final function (LP_Trend_Analysis) examines the engagement of other users within the pools, excluding the owners. It analyses transactions made by users to understand the general activity level and trends within each pool. This function helps to highlight user participation and can indicate the popularity of the LP based on user activity.
We further present pseudocode for our enhanced method by using machine learning (Algorithm 1).
Table IV: Summerized features for datasets.
@@
LP
$Address_{pool}$ Ethereum address of LPs $Address_{base}$ Ethereum address of base token in LPs
@@
DEXDEX DEX where the liquidity pool deployed NameName Paired token’s name
DEX activities
BlockBlock Block number that the activity occurred TimestampTimestamp Timestamp that the activity occured
HashHash Transaction hash of the activity CategoryCategory Activities (buy/sell/deposit/withdraw)
@@
Honeypot
$Tax_{buy}$ Buy tax of the token $Tax_{modify}$ Owner’s capability to change buy/sell
$Tax_{sell}$ Sell tax of the token BuyableBuyable User’s capability to buy tokens
SellAllSellAll User’s capability to sell all tokens BalanceChangeBalanceChange Owner’s capablity to change user’s balance
TradingCooldownTradingCooldown Time limit for tokens between activities TradingPausableTradingPausable Owner’s capbility of pausing trading
AntiWhaleAntiWhale token activities with high amounts
@@
OAF
$Owner_{dep}$ Total owner deposit activity $Owner_{buy}$ Total owner buy activity
@@
UAF
$User_{dep}$ Total users deposit activity $User_{with}$ Total users withdraw activity
@@
PF
$Owner_{i}$ Total owner investment (USD) $Owner_{u}$ Total owner unrealized return (USD)
@@
$ROwner_{roi}$ The ratio of ($Owner_{r}$-$Owner_{i}$) and $Owner_{i}$ $ROwner_{uonr}$ The ratio of $Owner_{u}$ and $Owner_{r}$
@@
LPF
AgeAge LP age (=d=d if the pool still active) IsAliveIsAlive Whether the pool is still active